(1 point) Find the Maclaurin series of the function f(x) = (2x2)e-9x. f(0) = Ź CT Determine the following coefficients:
Answer #1

solution 2 f(x) = e-gue - flors= Å Gnan ne (2 ) J 9= f te), ez e, C = Po TEMPAT $(x)= (43 41922) e*** +6)= 0 - 1 -9x (4-區) - REFE-555 国] = == =+ (ST-1 = = =) = = =)